Find the value of z such that 0.9282 of the area lies between −z and z. Round your answer to two decimal places.
Answer #1

Solution :

Using standard normal table,

P(-z < Z < z) = 0.9282

P(Z < z) - P(Z < -z) = 0.9282

2P(Z < z) - 1 = 0.9282

2P(Z < z) = 1 + 0.9282 = 1.9282

P(Z < z) = 1.9282 / 2 = 0.9641

P(Z < 1.80) = 0.9641

z = 1.80